# Measurement-based Fast Quantum State Stabilization with Deep Reinforcement Learning

### v1

@@ 5.2 GHZ State
where $n$ is the number of qubits. Its density matrix can be expressed as $\rho_{\rm GHZ}\triangleq|{\rm GHZ}\rangle\langle{\rm GHZ}|$. For the three-qubit GHZ state, we choose $|{\rm GHZ}\rangle=(1/\sqrt{2})({|000\rangle}+{|111\rangle})$, which gives the following density matrix:
@@
A degenerate observable $c$ is required according to the quantum state collapse after measurement [45, 8, 49]. The quantum state collapse states that the system in Eq. (1) will randomly converge to an eigenstate or eigenspace of $c$ without any control. Hence, we choose an observable in the following diagonal form:
@@
Due to the degenerate form of the observable $c$ in Eq. (19), the system may converge to other state in the corresponding eigenspace related to $\lambda_{d}$, two control channels $u_{1}$ and $u_{2}$ based on Lyapunov method have been applied in [45, 8, 49] to solve this problem. For subsequent performance comparisons, two control channels are also used in this paper.
@@
For all instances, in order to compare with the Lyapunov methods presented in [8], we choose the same system Hamiltonian as $H_{0}={\rm diag}[1,-1,-1,1,1,-1,-1,1]$, the target state as $\rho_{d}\triangleq\rho_{\rm GHZ}$ (18), and the observable $c$ as
@@
$={\rm diag}[3,1,-3,-1,-1,-3,1,3].$ (20)
The control Hamiltonians $H_{1}$ and $H_{2}$ are chosen as
$H_{1}$ $=I_{2}\otimes I_{2}\otimes\sigma_{x}+\sigma_{x}\otimes\sigma_{x}\otimes I_{2}$ (21)
$={\begin{bmatrix}0&1&0&0&0&0&-1&0\\ 1&0&0&0&0&0&0&-1\\ 0&0&0&1&-1&0&0&0\\ 0&0&1&0&0&-1&0&0\\ 0&0&-1&0&0&1&0&0\\ 0&0&0&-1&1&0&0&0\\ -1&0&0&0&0&0&0&1\\ 0&-1&0&0&0&0&1&0\end{bmatrix}},$
@@
$H_{2}$ $=\sigma_{x}\otimes I_{2}\otimes I_{2}+I_{2}\otimes\sigma_{x}\otimes\sigma_{x}$ (22)
$={\begin{bmatrix}1&0&0&1&0&0&0&0\\ 0&1&1&0&0&0&0&0\\ 0&1&1&0&0&0&0&0\\ 1&0&0&1&0&0&0&0\\ 0&0&0&0&-1&0&0&1\\ 0&0&0&0&0&-1&1&0\\ 0&0&0&0&0&1&-1&0\\ 0&0&0&0&1&0&0&-1\end{bmatrix}}.$
We test the trained DRL agent in various environments to evaluate its performance and robustness. The goal is to assess how well the agent generalizes its learned policies to different scenarios and how it copes with perturbations and variations in the environment. To achieve this, we expose the trained DRL agent to a set of diverse environments, each with unique characteristics and challenges. These environments are carefully designed to represent a wide range of scenarios and potential disturbances that the agent might encounter in real-world applications. During the testing phase, we measure the agent’s performance in terms of its ability to achieve the desired objectives and maintain stability in each environment. We also examine its response to changes in the measurement efficiency $\eta_{c}$ and time delay disturbances to assess its robustness and adaptability.
@@
5.2.1 Stabilization of the GHZ state under perfect case
@@
We initiate the testing phase to evaluate the ability of arbitrary initial states to stabilize to the target GHZ state within a specified time frame. As shown in Figure 4, we employ the comparative approach mentioned in Section 5.1, randomly selecting $50$ distinct initial states for control using the DRL agent and the Lyapunov method. The blue and orange lines correspond to the DRL method and the Lyapunov method, respectively. Under the guidance of the DRL agent, the control strategy improved the average evolution time by $22.38\%$ compared to the Lyapunov method (DRL: 33.39 a.u. vs. Lyapunov: 43.02 a.u.). This indicates that our DRL approach successfully stabilizes quantum states to the target GHZ state more rapidly.
@@
Figure 4: Evolution of the distance $D_{\rho_{t}}$ for $50$ random initial states stabilized to the target GHZ state under the control of the DRL agent (blue) and Lyapunov method (orange). The average stabilization time to the target under DRL control is $33.39$ a.u., while the Lyapunov method requires an average time of $43.02$ a.u.. (The stabilization time is defined as the time when the distance $D_{\rho_{t}}\leq 0.001$.) The light blue (orange) lines represent the average evolution trajectories for different initial states, and the dark blue (orange) line represents the average trajectory across all different initial states.
@@
In addition, we also explore the evolution of two specific initial states, denoted as $\rho_{0}^{1}={\rm diag}[0,1,0,0,0,0,0,0]$ and $\rho_{0}^{2}={\rm diag}[1,0,0,0,0,0,0,0]$, mentioned in [8] as examples. We repeat their stabilization $50$ times each to obtain averaged convergence curves that approximate the system’s evolution. Figure 5(a) and Figure 5(b) depict the evolution of these two distinct initial states. The blue curve represents the evolution controlled by the DRL agent, while the orange curve represents the evolution controlled by the Lyapunov method from [8]. It can be observed that the well-trained DRL agent not only achieves stable convergence to the target state but also showcases faster convergence compared to the Lyapunov method.
@@
Figure 5: Evolution of the distance $D_{\rho_{t}}$ to stabilize two particular initial states (a) $\rho_{0}^{1}={\rm diag}[0,1,0,0,0,0,0,0]$, (b) $\rho_{0}^{2}={\rm diag}[1,0,0,0,0,0,0,0]$ to the target GHZ state under the control of the DRL agent (blue) and the Lyapunov method (orange). Each trajectory represents the average of 50 different stabilization processes for these specific initial states.
@@
We randomly select a single trajectory under the control of a DRL agent with initial state $\rho_{0}^{1}$. The left subplot of Figure 6 illustrates the evolutionary trajectory with $D_{\rho_{t}}$, and the top images display the Wigner function for a harmonic mode of the system state at five different evolution times. In contrast, the subplot on the right serves as a reference plot for the target three-qubit GHZ state. A comparison reveals that the phase-space distribution of the system state gradually approaches the target state over time, and at $t=20$ a.u., the system state is identical to the target state.
@@
Figure 6: One specific (not averaged) evolutionary trajectory of the distance $D_{\rho_{t}}$ under the control of the DRL agent, starting from the initial state $\rho_{0}^{1}$. The top five images depict the Wigner function for a harmonic mode of the system at different moments during the evolution. The figure on the right shows the phase-space probability distributions of the standard three-qubit GHZ state, which serves as a reference.
@@
5.2.2 Stabilization of the GHZ state with imperfect measurement
We start from the measurement inefficiency. It’s important to note that we are using the agent trained under the assumption of “perfect case” to test the robustness of the agent. We assume a measurement efficiency $\eta_{c}=0.8$, which represents a relatively high efficiency achievable in current laboratory settings [50]. As shown in the Figure 7, $50$ randomly selected initial states are successfully stabilized to the target GHZ state under the control of the agent. Therefore, our trained DRL agent has shown good robustness in terms of measurement inefficiency.
@@ 5.2.2 Stabilization of the GHZ state with imperfect measurement
Figure 7: With a measurement efficiency $\eta_{c}=0.8$, under the control of the DRL agent, 50 random initial states stably evolve to the target GHZ state. The light blue lines represent the average evolution trajectories for different initial states, and the dark blue line represents the average trajectory across all different initial states.
5.2.3 Stabilization of the GHZ state with time delay
We then consider the issue of time delay in feedback processes. Especially in the rapid evolution of quantum systems, the time required for traditional computers to solve the SME (1) is often non-negligible. Therefore, it is worth considering incorporating a fixed compensation for time during the testing process of the agent. For instance, assuming a time delay of $\mathcal{t}=0.05$ a.u., starting from $t=0$ a.u. until $t=0.05$ a.u., the agent receives only the initial state $\rho_{0}$ as input and provides signals to control the system’s evolution based on the input state. The agent receives the state $\rho_{1}$ at $t=0.05$ a.u., and so on, where at each step, the agent receives the state $\rho_{t-0.05}$, which is the state prior to $\mathcal{t}=0.05$ a.u.. As shown in the Figure 8, by selecting $50$ different initial states arbitrarily, it can be observed that a well-trained agent still exhibits excellent performance in dealing with time delays, although longer convergence time is required than in the perfect environment in Figure 4.

### v2

@@ V-B GHZ State
where $n$ is the number of qubits. Its density matrix can be expressed as $\rho_{\rm{GHZ}}\rm Triangleq|{\rm{GHZ}}\rangle\langle{\rm{GHZ}}|$. For the three-qubit GHZ state, we choose $|{\rm{GHZ}}\rangle=(1/\sqrt{2})({|000\rangle}+{|111\rangle})$, which gives the following density matrix:
@@
A degenerate observable $c$ is required according to the quantum state collapse after measurement [36, 8, 40]. The quantum state collapse states that the system in (1) will randomly converge to an eigenstate or eigenspace of $c$ without any control. Hence, we choose an observable in the following diagonal form:
@@
Due to the degenerate form of the observable $c$ in (14), the system may converge to other state in the corresponding eigenspace related to $\lambda_{d}$, two control channels $H_{1}$ and $H_{2}$ based on Lyapunov method have been applied in [36, 8, 40] to solve this problem. For subsequent performance comparisons, two control channels are also used in this paper.
@@
For all instances, in order to compare with the Lyapunov methods presented in [8], we choose the same system Hamiltonian as $H_{0}=\text{diag}[1,-1,-1,1,1,-1,-1,1]$. The diagonal form of which indicates that the eigenvalues correspond to the energy levels of the system in the computational basis $|q_{1}q_{2}q_{3}\rangle$ (where $q_{i}\in\{0,1\}$ for three qubits). The target state is $\rho_{d}\rm Triangleq\rho_{\rm{GHZ}}$ (13), and the observable $c$ is chosen to be:
@@
$=\text{diag}[3,1,-3,-1,-1,-3,1,3],$ (15)
to measure correlations between the z-components of different pairs of qubits in a three-qubit system and also ensure that the target state is an eigenstate. The control Hamiltonians $H_{1}$ and $H_{2}$ are chosen as
$H_{1}=I_{2}\otimes I_{2}\otimes\sigma_{x}+\sigma_{x}\otimes\sigma_{x}\otimes I_{2},$ (16)
@@
$H_{2}=\sigma_{x}\otimes I_{2}\otimes I_{2}+I_{2}\otimes\sigma_{x}\otimes\sigma_{x}.$ (17)
Similar to the two-qubit systems, two control channels represented by $H_{1}$ and $H_{2}$ together with their strengths $u_{1}$ and $u_{2}$ provide control over the system’s dynamics. The forms of control Hamiltonian represent how the control action is applied to the three-qubit system. For example, the form of $H_{1}$ generally represents an independent $x$-axis control to the third qubit and a correlated $x$-axis interaction between the first and the second qubits.
We test the trained DRL agent in various environments to evaluate its performance and robustness. The goal is to assess how well the agent generalizes its learned policies to different scenarios and how it copes with perturbations and variations in the environment. To achieve this, we expose the trained DRL agent to a set of diverse environments, each with unique characteristics and challenges. These environments are carefully designed to represent a wide range of scenarios and potential disturbances that the agent might encounter in real-world applications. During the testing phase, we measure the agent’s performance in terms of its ability to achieve the desired objectives and maintain stability in each environment. In particular, we examine its response to changes in the measurement efficiency $\eta_{c}$ and time delay disturbances to assess its robustness and adaptability.
@@
V-B1 Stabilization of the GHZ state under perfect case
@@
We initiate the testing phase to evaluate the ability of arbitrary initial states to stabilize to the target GHZ state within a specified time frame. As shown in Fig. 5, we employ the comparative approach mentioned in Section V-A, randomly selecting $50$ distinct initial states for control using the DRL agent and the Lyapunov method. The blue and orange lines correspond to the DRL method and the Lyapunov method, respectively. The average evolution time of 2500 trajectories guided by the DRL agent is $16\%$ shorter than that of the Lyapunov method (DRL: 10.41 a.u. vs. Lyapunov: 12.33 a.u.). This indicates that our DRL approach successfully stabilizes quantum states to the target GHZ state more rapidly.
@@
Figure 5: Evolution of the distance $D_{\rho_{t}}$ for $50$ random initial states stabilized to the target GHZ state under the control of the DRL agent (blue) and Lyapunov method (orange). The average stabilization time to the target under DRL control is $10.41$ a.u., while the Lyapunov method requires an average time of $12.33$ a.u.. (The stabilization time is defined as the time when the distance $D_{\rho_{t}}\leq 0.001$.) The light blue (orange) lines represent the average evolution trajectories for different initial states, and the dark blue (orange) line represents the average trajectory across all different initial states.
@@
In addition, we also explore the evolution of two specific initial states, denoted as $\rho_{0}^{1}=\text{diag}[0,1,0,0,0,0,0,0]$ and $\rho_{0}^{2}=\text{diag}[1,0,0,0,0,0,0,0]$, mentioned in [8] as examples. We repeat their stabilization $50$ times each to obtain averaged convergence curves that approximate the system’s evolution. Fig. 6(a) and Fig. 6(b) depict the evolution of these two distinct initial states. The blue curve represents the evolution controlled by the DRL agent, while the orange curve represents the evolution controlled by the Lyapunov method from [8]. It can be observed that the well-trained DRL agent not only achieves stable convergence to the target state but also showcases faster convergence compared to the Lyapunov method.
@@
Figure 6: Evolution of the distance $D_{\rho_{t}}$ to stabilize two particular initial states (a) $\rho_{0}^{1}=\text{diag}[0,1,0,0,0,0,0,0]$, (b) $\rho_{0}^{2}=\text{diag}[1,0,0,0,0,0,0,0]$ to the target GHZ state under the control of the DRL agent (blue) and the Lyapunov method (orange). Each trajectory represents the average of 50 different stabilization processes for these specific initial states.
@@
We randomly select a single trajectory under the control of a DRL agent with initial state $\rho_{0}^{1}$. The left subplot of Fig. 7 illustrates the evolutionary trajectory with $D_{\rho_{t}}$, and the top images display the Wigner function of the system state at five different evolution times. In contrast, the subplot on the right serves as a reference plot for the target three-qubit GHZ state. A comparison reveals that the phase-space distribution of the system state gradually approaches the target state over time, and at $t=20$ a.u., the system state is identical to the target state.
@@
Figure 7: One specific (not averaged) evolutionary trajectory of the distance $D_{\rho_{t}}$ under the control of the DRL agent, starting from the initial state $\rho_{0}^{1}$. The top five images depict the Wigner function of the system state at different moments during the evolution. The figure on the right shows the phase-space probability distributions of the standard three-qubit GHZ state, which serves as a reference.
@@
V-B2 Stabilization of the GHZ state with imperfect measurement
We first investigate the impact of reduced measurement efficiency, focusing on the robustness of an agent trained under the assumption of “perfect case”. In this test, we consider a measurement efficiency of $\eta_{c}=0.8$, which represents a relatively high level achievable in current laboratory environments [41]. As shown in Fig. 8, both the DRL-based agent and the Lyapunov-based method successfully stabilize $50$ randomly selected initial states to the target GHZ state. Moreover, the DRL agent demonstrates slightly superior performance. These results highlight that the DRL agent, trained under ideal conditions, retains significant robustness even in the presence of reduced measurement efficiency.
@@ V-B2 Stabilization of the GHZ state with imperfect measurement
Figure 8: With a measurement efficiency $\eta_{c}=0.8$, under the control of the DRL agent and the Lyapunov-based method, $50$ random initial states stably evolve to the target GHZ state. The light lines represent the average evolution trajectories for different initial states, and the dark blue line represents the average trajectory across all different initial states.
V-B3 Stabilization of the GHZ state with time delay
We evaluate the performance of the trained agent under the presence of time delays in the feedback process. In rapidly evolving quantum systems, the time required for traditional computers to solve the SME (1) is often non-negligible. To account for this, we incorporate fixed time compensation during agent testing. For example, assuming a time delay of $\mathcal{t}=0.05$ a.u., the agent only receives the initial state $\rho_{0}$ as input from $t=0$ a.u. to $t=0.05$ a.u., and generates control signals based on this state to guide the system’s evolution. At $t=0.051$ a.u., the agent receives the state $\rho_{1}$, and subsequently at each step, it processes the state $\rho_{t-0.05}$, which corresponds to the state from $\mathcal{t}=0.05$ a.u. earlier. As illustrated in Figure 9, using $50$ randomly selected initial states, we observe that both the trained DRL agent and the Lyapunov-based method handle time delays effectively. Moreover, the DRL agent demonstrates superior performance compared to the Lyapunov-based approach.